OpenStudy (anonymous):

f (x)=x^2-6 @(-5,19)

OpenStudy (gorv):

f(x) is nothing but the value of y y=f(x)=x^2-6 now plug in the value of x y=(-5)^2-6 value will come out 19 as given in the Q so this point lie on this curve